# Bounding field excursions along null geodesics with applications to cosmology

### v1

@@ 1 Introduction
Quantum field theories often possess a moduli space of vacua parameterized by the expectation values of scalar fields. When the expectation values vary coherently in spacetime, gravity couples and responds to these variations. Such variations are not merely an academic possibility: in the early universe, large-scale gradients in scalar field expectation values may have driven inflationary expansion [1, 2], leaving imprints in the cosmic microwave background, while in extreme astrophysical environments, such as near-extremal black holes [3], coherent moduli gradients can persist over macroscopic distances. [^1] Israel’s third law implies that a (near-)extremal black hole cannot be produced by any finite, quasi-stationary physical process, thereby naively ruling out the astrophysical formation of such objects [4]. However, arguments for Israel’s third law omit quantum effects; in particular, Hawking evaporation can drive a magnetically charged black hole arbitrarily close to extremality on a timescale shorter than the age of the universe [5]. Even modest departures from uniformity can deflect light rays and shift horizons in spacetime. Taken together with the ongoing debate over whether distinct vacua belong to a single quantum theory [6, 7, 8, 9, 10], these effects underscore that spatially varying vacua are a fundamental component of any realistic description of gravity–matter systems.
@@
In this paper, we study bounds on field excursions in localized regions due to gravitational backreaction. An early indication of bounds on field excursions in localized regions appeared in Ref. [11], which showed that trans-Planckian field configurations can be constructed within the Newtonian approximation, but only if the spatial extent of the experiment grows exponentially as a function of the excursion distance in the theory moduli space. Subsequent studies explored field variations in more phenomenological settings, including cosmic strings and bubbles of nothing [12, 13, 14]. More recently, Ref. [15] demonstrated that regions deep in the vacuum manifold can be accessed using extremal black holes, whose entropy again grows exponentially with the field distance. [^2] See also related work in Refs. [16, 17, 18, 19].
@@
Building on these results, we derive a quantitative bound, the field excursion bound (FEB), on field excursions along null geodesics, in terms of the expansion parameter $\theta$. The core intuition is that coherent variations in the expectation value of scalar fields induce gravitational focusing of light rays, and the expansion parameter $\theta$ quantifies the extent of this focusing. The Raychaudhuri equation [20, 21], the scalar action, and the null energy condition (NEC) for all other fields enable us to derive the FEB. We verify the FEB in spacetimes containing a timelike naked singularity, where it is saturated.
@@
Our primary motivation for introducing the FEB comes from anthropic considerations in inflationary modes of early-universe cosmology. In anthropic models of inflation, where the existence of observers guides vacuum selection, it is often assumed that scalar fields must traverse large distances in the theory moduli space to reach vacua with suitable low-energy properties. In cosmological settings, the FEB implies that the extent of large field excursions is linearly upper-bounded by the number of e-folds, independent of the underlying cosmological model. This leads to a striking implication: if an inflationary multiverse is indeed responsible for the observed fine-tuning of our universe, then the inflationary epoch should have lasted many e-folds. This prediction stands in significant tension with the Hartle-Hawking wavefunction [22], which favors initial conditions with minimal inflation [23].
@@
We also propose a potential generalization of the FEB to semiclassical spacetimes that violate the NEC. Our proposal relies on a strengthened Quantum Focusing Condition (QFC) [24]. Although we do not prove the strengthened QFC, we note that it leads to some desirable properties of the spacetime beyond a quantum generalization of the FEB. Although the quantum FEB and strengthened QFC have no immediately evident phenomenological applications, they are nonetheless of significant formal interest and may prove relevant in future contexts.
@@ 2 The field excursion bound
To state the FEB, we first briefly review the geometry of null congruences. [^3] See Appendix F of Ref. [25] for a detailed discussion. Consider a family of null geodesics with an affinely parameterized, tangent vector field $k^{\mu}$. Introduce an auxiliary null vector $l^{\mu}$ such that both $k^{\mu}$ and $l^{\mu}$ are orthogonal to a codimension-two spatial hypersurface $\sigma$, whose evolution along the congruence we wish to track. We impose the conditions
@@
The null Raychaudhuri equation [20, 21] implies that
@@ 3 Example: naked time-like singularity
Field excursions can be arbitrarily large in the presence of time-like naked singularities and thereby provide a nice cross-check of the FEB. We consider the class of solutions given in Ref. [26],444and further studied in Refs. [11]. which considered an algorithm for constructing a solution with a non-trivial scalar field profile given a vacuum solution where the scalar field profile is trivial. The scalar is taken to be minimally coupled to gravity. Taking the Schwarschild solution as input, the algorithm outputs the metric
@@
Note that the expansion parameter diverges at $r=2m$ unless $\beta=1$ [25]. To cross-check the FEB, we want to check the FEB at every point along a geodesic. We therefore consider the infinitesimal change in $\log(\theta)$ relative to $\phi$:
@@ 4 Example: large-field inflation
In this section, we consider the FEB in the context of inflation. Inflation is currently the leading model of early universe cosmology as it explains the horizon and flatness problems [27, 28, 29]. [^5] See Ref. [30] for a nice review. The general idea is that there is a field called the inflaton that transversed a tall, flat potential. Due to the high potential energy of the inflation during this period, the universe was rapidly expanding and de-Sitter-like. Eventually, the inflaton fell into a well in $V(\phi)$, which led to reheating and the near-flat cosmology we are familiar with.
@@ 4.2 Less symmetrical inflation
Here, the first inequality holds by construction, because $s_{i}$ has $\mathrm{tr}\,k=0$, while $\kappa_{i}$ has $\mathrm{tr}\,k>0$ (for the outward normal $n^{a}$, which points in the opposite direction to $k^{a}$). The second inequality holds by virtue of the MCA, because $\gamma_{i}$ is marginally bounded by $\kappa_{i}$, meaning that $\gamma_{i}$ (unless it coincides with $\kappa_{i}$ in a neighborhood of the point $q$) can only deviate from $\kappa_{i}$ in an inward direction, not an outward direction (see the discussion of this geometric fact in [32, 33, 34]). And the third inequality has already been stated above.
@@ 4.3 Applications to phenomenology
As an example of how this result might affect phenomenology, let us suppose we are trying to use inflation to solve the fine-tuning problems of theoretical particle physics, by postulating a large landscape of $n_{v}$ possible vacua. [^14] Here, we leave aside the important philosophical question about whether a multiverse would count as a good explanation for fine-tuning (see e.g. [35, 36] and citations therein). We confine our analysis to the physics question of whether there is enough time to reach the requisite number of vacua. To illustrate the point in the starkest possible terms, let us suppose that $\ell_{\text{string}}\sim\ell_{\text{SUSY}}\sim\ell_{\text{KK}}\sim\ell_{\text{planck}}$ (assuming that, respectively, string theory, supersymmetry, and/or Kaluza-Klein dimensions exist), and that at energy scales below $\ell_{\text{planck}}^{-1}$, physics is well-described by a traditional renormalizable Wilsonian QFT paradigm in $D=4$ spacetime dimensions. (In other words, we take away all other resources that would be helpful for reducing the amount of fine-tuning, and try to solve it with a landscape alone.) Between the cosmological constant (tuned to $\sim\!10^{-120}$) and the Higgs mass squared (tuned to $\sim\!10^{-30}$), we would need $n_{v}\gtrsim 10^{150}$ distinct vacua to expect to find at least one vacuum that is as fine-tuned as our universe. [^15] If we instead postulate low energy supersymmetry ($\ell_{\text{SUSY}}\sim 100\text{ TeV}$), the amount of fine-tuning required drops to around $10^{-60}$, and it is also permissible for the scalar potentials to be less generic. But this requires restricting to supersymmetric regions of moduli space, and it is not totally clear whether such vacua are more numerous in the string landscape than vacua with Planck scale (or no) supersymmetry [37], even though the latter require more fine-tuning.
@@
where we ignore $\phi^{5}$ and higher terms as they are irrelevant in $D=4$. In a generic situation with no symmetry, we would expect $\lambda^{(4)}_{AB}$ to take a generic positive quartic form. [^16] There might be ways of cleverly engineering additional vacua using e.g. non-polynomial terms (from radiative corrections) in effective quantum potentials, but we assume this is unlikely to produce an enormously greater number of vacua. The equation $V^{\prime}=0$ is thus a system of $n_{s}$ cubic equations. For a generic quartic potential, Bézout’s theorem [^17] Bézout’s theorem asserts that a system of $n$ polynomial equations in $n$ variables has exactly $\prod_{i=1}^{n}\deg(f_{i})$ solutions (counted with multiplicity, and including complex solutions as well as projective solutions “at infinity”), except in certain nongeneric cases, where the number of solutions is instead infinite. implies that there are exactly $3^{n_{s}}$ distinct critical points, which might, however, be at either real or complex values of the $\phi^{A}$ coordinates. In our case we are interested only in the critical points which are i) real, ii) stable, and iii) within the regime of validity of the potential (46), so we get a (not always tight) upper bound of $3^{n_{s}}$ stable vacua within the ball $B$, i.e. that are $\ll 1$ Planck-excursion from the original vacuum. [^18] A simple example of a potential with $3^{n_{s}}$ real critical points is $V=\sum_{A}\frac{1}{4}(\phi^{A})^{4}-\frac{1}{2}(\phi^{A})^{2}$. In this case, only $2^{n_{s}}$ of them are stable. Surprisingly, this is not the actual upper bound! In the case $n_{s}=2$, [38] proved that there cannot be more than five stable critical points, while the functions described in [39] and [40] show that this can, in fact, be achieved. Of course, there may also be many other planck-sized regions of moduli space that contain no vacua.
@@
The central point is that if an inflationary multiverse truly accounts for fine-tuning, then (unless there are a rather large number of scalar fields) the inflationary epoch must go back much earlier than the current observational lower bound. This expectation contrasts sharply with the Hartle–Hawking wavefunction proposal [22]. Although the proposal appears to give reasonable predictions for the high-energy modes of the quantum fluctuations [41, 42, 43, 44, 45], it favors an inflationary period with far fewer e-folds than even observational bounds permit [23]. In fact, if we do not condition on inflation happening at all, the Hartle-Hawking proposal predicts that an even more dominant contribution to the saddle should be a big empty de Sitter universe with the present-day value of $\Lambda$, dramatically contradicting observation.
These issues could potentially be remedied by considering another variant of the no-boundary proposal, for example the Vilenkin tunnelling proposal [^19] This model was originally proposed in [46], for more recent discussion see Vilenkin and Yamada [47, 48] and citations therein. in which the universe tunnels from zero size ($a=0$) along a Lorentzian contour. Such models typically make the opposite prediction from the Hartle-Hawking model, that the universe should start out as small as possible, and thus at a high value of the inflationary potential. It was argued in [49] that this model gives an unnormalizable (inverse Gaussian) prediction for the quantum fluctuations, but Vilenkin and Yamada claim that this issue can be resolved by adding an appropriate boundary term to the action; see also [50] for a similar idea.
Thus, if the initial condition was not Hartle-Hawking, it is more plausible that the universe started at a point in the landscape from which a large amount of inflation was possible. It is also possible for inflation never to end (see e.g. [51] for a review of this “eternal inflation”), but as such scenarios typically involve quantum fluctuations, the above conclusions will not necessarily be applicable unless the FEB can be generalized to quantum situations. This will be the subject of the next section.
@@ 5 Quantum effects
We now comment on a possible quantum generalization. As stated above Eq. (8), we assumed the NEC in this derivation. However, the NEC can be violated by quantum effects [52] such as Hawking radiation. We need to consider a quantum generalization of the NEC, which will involve accounting for entropy variation. We will keep Newton’s constant, $G_{\textrm{N}}$, explicit in this section because its renormalization will play an important role.
@@
associated with a given $\sigma_{\lambda}(y_{\perp})$. $S_{\text{gen}}$ generalizes the notion of area when quantum effects are important. Rather than studying how the entropy and area evolve as a function of $\lambda$, we treat the area and entropy as functionals of $\sigma_{\lambda}(y_{\perp})$ and consider functional derivatives localized in a small area $\mathcal{A}$ around $y_{1,\perp}$; see Fig. 3. Using the same notation as Ref. [24], we use the shorthand
@@
$\frac{d\theta}{d\lambda}\leq\frac{-1}{D-2}\theta^{2}$ (55)
@@
motivated by the idea that quantum focusing inequality should have a definite-sign quadratic term just as the classical Raychaudhuri Eq. does. Then we consider the contribution of a coherently varying field along the light ray. We refer to Eq. (56) as the SQFC, because the original quantum focusing conjecture (QFC) only imposed that the left-hand side was negative [24].
@@
In the original QFC, the entropy suffers from UV divergences, but these can be absorbed into the renormalization of Newton’s constant, so $\Theta$ is well-defined [53, 24]. More concretely, the leading order area-law renormalization leads to the replacements
@@
for some constant $b$, where $G_{\textrm{N}}$ is Newton’s constant. The shift in Eq. (58) leaves $S_{\textrm{gen}}$ (52) invariant. In contrast, a shift in Newton’s constant due to renormalization changes the relative sizes of the two terms in Eq. (57) due to the explicit factor of $G_{\textrm{N}}$. The meaning of the bound thus requires picking some choice of $G_{\textrm{N}}$, which is a nontrivial decision if $G_{\textrm{N}}$ is subject to RG flow. As a conservative choice, we propose that whenever the SQFC is applied to an EFT calculation valid between some minimum energy scale $E_{\mathrm{min}}$ and maximum energy scale $E_{\mathrm{max}}$, that (i) an effective value of $G_{N}(E)$ should be defined at each energy $E_{\text{min}}<E<E_{\text{max}}$ using the area-law formula, and that (ii) the SQFC holds for whichever of these values of $G_{N}(E)$ is the largest (i.e. the least stringent choice in (58)).
@@
and an “area law” scheme for $G_{\text{N}}$ will set $c_{1}=0$, so that the area-law part of the generalized entropy is entirely contained in the $A/4G_{\text{N}}$ term. A more complicated story occurs if the matter sector flows from a UV CFT to an IR CFT. In this case, Eq. (59) only holds asymptotically in either the $r\to 0$ limit (the UV CFT) or the $r\to\infty$ limit (the IR CFT). It has been proven that $S^{\prime\prime}(r)<0$ [54], so it is not possible to have $c_{1}=0$ in both the UV and the IR. In fact, $G_{\text{UV}}<G_{\text{IR}}$ (i.e. in this area-law scheme gravity is anti-screening). In the case of the future lightcone of a point $p$ in the 2+1 Minkowski vacuum, we have checked that the SQFC holds when defined in terms of $G_{\text{UV}}$, by setting $c_{1}=0$ in the UV. (This implies that $S^{\prime}(r)\leq 0$ for all $r>0$, and thus ensures that $\theta$ and $S^{\prime}$ have the opposite signs.)
@@
In higher dimensions, it might be necessary to also think carefully about subleading divergences, beyond the area law, which are associated with higher curvature terms in the effective gravitational action (cf. [24] and citations therein).
@@
The strengthened QFC implies some desirable properties about spacetime, such as quantum generalizations of the Penrose singularity theorem [55, 56, 57]. It was previously argued by one of us that the Generalized Second Law (GSL), which is implied by the QFC, can be used to prove the incompleteness of null geodesics from the existence of a “quantum trapped surface” $\cal T$ with $\Theta\leq 0$ everywhere [32]. [^20] Together with the auxiliary assumptions of global hyperbolicity and spatial noncompactness, which are also used in the classical Penrose theorem. The SQFC would allow one to strengthen this result by deriving that the termination should occur before an affine time $\Delta\lambda\leq-(D-2)/\Theta[\mathcal{T}]$, analogously to the classical result.
@@
We further note that the SQFC is similar to a bound proposed by Ben-Dayan [58]. Although the details of Ben-Dayan’s inequality vary from our own, it is similarly obtained by adding to the QFC a term which vanishes when $\Theta=0$. All such bounds will imply the restricted QNEC, that $d\Theta/d\lambda\leq 0$ whenever $\Theta=0$, which has been proven in the context of the brane-worlds scenario [59].
@@
We now turn back to bounding field excursions. The core intuition behind the original FEB’s derivation from the NEC is that, if the other stress–energy contributions violated the NEC, one could simply eliminate the field excursion and reveal a NEC violation. We adopt a similar line of reasoning here. We assume that the expectation values of the scalar field and the area operator obey Einstein’s equations. To derive a quantum FEB (QFEB), we consider an original state where the scalar expectation value $\langle\phi\rangle$ changes along the null ray. We then act on this state with an operator that removes the coherent variation of $\phi$ within a local region and apply Eq. (56) to this state. Thus, in the new state $\langle\phi\rangle=0$. (As the local region can be small, the required shift in $\Delta\langle\phi\rangle$ can be made arbitrarily small, and thus should not significantly affect the behavior of the matter quantum state through direct couplings.) In order for Eq. (56) to hold for the new state, the following bound must hold on the original state with coherent field variation:

### v2

@@ 1 Introduction
Quantum field theories often possess a moduli space of vacua parameterized by the expectation values of scalar fields. When the expectation values vary coherently in spacetime, gravity couples and responds to these variations. Such variations are not merely an academic possibility: in the early universe, large-scale gradients in scalar field expectation values may have played a crucial role in inflation [1983veu..conf..251S, Vilenkin:1983xq], leaving imprints in the cosmic microwave background, while in extreme astrophysical environments, such as near-extremal black holes [Ferrara:1995ih], coherent moduli gradients can persist over macroscopic distances. [^1] Arguments for Israel’s third law suggest that a (near-)extremal black hole cannot be produced by any finite, quasi-stationary physical process, thereby naively ruling out the astrophysical formation of such objects [Israel:1986gqz]. However, arguments for Israel’s third law omit quantum effects; for example, Hawking evaporation can drive a magnetically charged black hole arbitrarily close to extremality on a timescale shorter than the age of the universe [Maldacena:2020skw]. See also Ref. [Kehle:2022uvc] for possible counterexamples to Israel’s third law in classical contexts. Even modest departures from uniformity can deflect light rays and shift horizons in spacetime. Taken together with the ongoing debate over whether distinct vacua in a gravitational theory belong to a single quantum theory [Banks:2000zy, Banks:2003vp, Banks:2019oiz, Banks:2025nfe, Sen:2025bmj], these effects underscore that spatially varying vacua are a fundamental component of any realistic description of gravity–matter systems.
@@
In this paper, we study bounds on field excursions in localized regions due to gravitational backreaction. An early indication of bounds on field excursions in localized regions appeared in Ref. [Nicolis:2008wh], which showed that trans-Planckian field configurations can be constructed within the Newtonian approximation, but only if the spatial extent of the experiment grows exponentially as a function of the excursion distance in the theory moduli space. Subsequent studies explored field variations in more phenomenological settings, including cosmic strings and bubbles of nothing [Dolan:2017vmn, Draper:2019zbb, Draper:2019utz]. More recently, Ref. [Delgado:2022dkz] demonstrated that regions deep in the vacuum manifold can be accessed using extremal black holes, whose entropy again grows exponentially with the field distance. [^2] See also related work in Refs. [Klaewer:2016kiy, Chaudhary:2020yyv, Sen:2025ljz, Sen:2025oeq].
@@
Building on these results, we derive a quantitative bound, the field excursion bound (FEB), on field excursions along null geodesics, in terms of the expansion parameter $\theta$. The core intuition is that coherent variations in the expectation value of scalar fields induce gravitational focusing of light rays, and the expansion parameter $\theta$ quantifies the extent of this focusing. The Raychaudhuri equation [Raychaudhuri:1953yv, Sachs:1961zz], the scalar action, and the null energy condition (NEC) for all other fields enable us to derive the FEB. We verify the FEB in a spacetime containing a timelike naked singularity, where it is saturated.
@@
Our primary motivation for introducing the FEB comes from anthropic considerations in inflationary models of early-universe cosmology. In anthropic models of inflation, where the existence of observers guides vacuum selection, it is often assumed that scalar fields must traverse large distances in the theory moduli space to reach vacua with suitable low-energy properties. In cosmological settings, the FEB implies that the extent of large field excursions is linearly upper-bounded by the number of e-folds, independent of the underlying cosmological model. This leads to a striking implication: if an inflationary multiverse is indeed responsible for the observed fine-tuning of our universe, then the inflationary epoch should have lasted many e-folds. This prediction stands in significant tension with the Hartle-Hawking wavefunction [Hartle:1983ai], which favors initial conditions with minimal inflation [Maldacena:2024uhs].
@@
We also propose a potential generalization of the FEB to semiclassical spacetimes that violate the NEC. Our proposal relies on a strengthened Quantum Focusing Condition (QFC) [Bousso:2015mna]. Although we do not prove the strengthened QFC, we note that it leads to some desirable properties of the spacetime beyond a quantum generalization of the FEB. Although the quantum FEB and strengthened QFC have no immediately evident phenomenological applications, they are nonetheless of significant formal interest and may prove relevant in future contexts.
@@ 2 The field excursion bound
To state the FEB, we first briefly review the geometry of null congruences. [^3] See Appendix F of Ref. [Carroll:2004st] for a detailed discussion. Consider a family of null geodesics with an affinely parameterized, tangent vector field $k^{\mu}$. Introduce an auxiliary null vector $l^{\mu}$ such that both $k^{\mu}$ and $l^{\mu}$ are orthogonal to a codimension-two spatial hypersurface $\sigma$, whose evolution along the congruence we wish to track. We impose the conditions
@@
The null Raychaudhuri equation [Raychaudhuri:1953yv, Sachs:1961zz] implies that
@@ 3 Example: naked time-like singularity
Field excursions can be arbitrarily large in the presence of time-like naked singularities and thereby provide a nice cross-check of the FEB. We consider the class of solutions given in Ref. [Buchdahl:1959nk],444and further studied in Refs. [Nicolis:2008wh]. which considered an algorithm for constructing a solution with a non-trivial scalar field profile given a vacuum solution where the scalar field profile is trivial. The scalar is taken to be minimally coupled to gravity. Taking the Schwarschild solution as input, the algorithm outputs the metric
@@
Note that the expansion parameter diverges at $r=2m$ unless $\beta=1$ [Carroll:2004st]. To cross-check the FEB, we want to check the FEB at every point along a geodesic. We therefore consider the infinitesimal change in $\log(\theta)$ relative to $\phi$:
@@ 4 Example: large-field inflation
In this section, we consider the FEB in the context of inflation. Inflation is currently the leading model of early universe cosmology as it explains the horizon and flatness problems [Guth:1980zm, Linde:1981mu, Albrecht:1982wi]. [^5] See Ref. [Baumann:2009ds] for a nice review. The general idea is that there is a field called the inflaton that transversed a tall, flat potential, $V(\phi)$. Due to the high potential energy of the inflation during this period, the universe was rapidly expanding and de-Sitter-like. Eventually, the inflaton fell into a well in $V(\phi)$, which led to reheating and the near-flat cosmology we are familiar with.
@@ 4.2 Less symmetrical inflation
Here, the first inequality holds by construction, because $s_{i}$ has $\mathrm{tr}\,k=0$, while $\kappa_{i}$ has $\mathrm{tr}\,k>0$ (for the outward normal $n^{a}$, which points in the opposite direction to $k^{a}$). The second inequality holds by virtue of the MCA, because $\gamma_{i}$ is marginally bounded by $\kappa_{i}$, meaning that $\gamma_{i}$ (unless it coincides with $\kappa_{i}$ in a neighborhood of the point $q$) can only deviate from $\kappa_{i}$ in an inward direction, not an outward direction (see the discussion of this geometric fact in [Wall:2010jtc, Wall:2012uf, Engelhardt:2013tra]). And the third inequality has already been stated above.
@@ 4.3 Applications to phenomenology
As an example of how this result might affect phenomenology, let us suppose we are trying to use inflation to solve the fine-tuning problems of theoretical particle physics, by postulating a large landscape of $n_{v}$ possible vacua. [^14] Here, we leave aside the important philosophical question about whether a multiverse would count as a good explanation for fine-tuning (see e.g. [manson2022cosmic, isaacs2022multiple] and citations therein). We confine our analysis to the physics question of whether there is enough time to reach the requisite number of vacua. To illustrate the point in the starkest possible terms, let us suppose that $\ell_{\text{string}}\sim\ell_{\text{SUSY}}\sim\ell_{\text{KK}}\sim\ell_{\text{planck}}$ (assuming that, respectively, string theory, supersymmetry, and/or Kaluza-Klein dimensions exist), and that at energy scales below $\ell_{\text{planck}}^{-1}$, physics is well-described by a traditional renormalizable Wilsonian QFT paradigm in $D=4$ spacetime dimensions. (In other words, we take away all other resources that would be helpful for reducing the amount of fine-tuning, and try to solve it with a landscape alone.) Between the cosmological constant (tuned to $\sim\!10^{-120}$) and the Higgs mass squared (tuned to $\sim\!10^{-30}$), we would need $n_{v}\gtrsim 10^{150}$ distinct vacua to expect to find at least one vacuum that is as fine-tuned as our universe. [^15] If we instead postulate low energy supersymmetry ($\ell_{\text{SUSY}}\sim 100\text{ TeV}$), the amount of fine-tuning required drops to around $10^{-60}$, and it is also permissible for the scalar potentials to be less generic. But this requires restricting to supersymmetric regions of moduli space, and it is not totally clear whether such vacua are more numerous in the string landscape than vacua with Planck scale (or no) supersymmetry [Susskind:2004uv], even though the latter require more fine-tuning.
@@
where we ignore $\phi^{5}$ and higher terms as they are irrelevant in $D=4$. In a generic situation with no symmetry, we would expect $\lambda^{(4)}_{AB}$ to take a generic positive quartic form. [^16] There might be ways of cleverly engineering additional vacua using e.g. non-polynomial terms (from radiative corrections) in effective quantum potentials, but we assume this is unlikely to produce an enormously greater number of vacua. The equation $V^{\prime}=0$ is thus a system of $n_{s}$ cubic equations. For a generic quartic potential, Bézout’s theorem [^17] Bézout’s theorem asserts that a system of $n$ polynomial equations in $n$ variables has exactly $\prod_{i=1}^{n}\deg(f_{i})$ solutions (counted with multiplicity, and including complex solutions as well as projective solutions “at infinity”), except in certain nongeneric cases, where the number of solutions is instead infinite. implies that there are exactly $3^{n_{s}}$ distinct critical points, which might, however, be at either real or complex values of the $\phi^{A}$ coordinates. In our case we are interested only in the critical points which are i) real, ii) stable, and iii) within the regime of validity of the potential (46), so we get a (not always tight) upper bound of $3^{n_{s}}$ stable vacua within the ball $B$, i.e. that are $\ll 1$ Planck-excursion from the original vacuum. [^18] A simple example of a potential with $3^{n_{s}}$ real critical points is $V=\sum_{A}\frac{1}{4}(\phi^{A})^{4}-\frac{1}{2}(\phi^{A})^{2}$. In this case, only $2^{n_{s}}$ of them are stable. Surprisingly, this is not the actual upper bound! In the case $n_{s}=2$, [durfee1993counting] proved that there cannot be more than five stable critical points, while the functions described in [MOexample] and [MSEexample] show that this can, in fact, be achieved. Of course, there may also be many other planck-sized regions of moduli space that contain no vacua.
@@
The central point is that if an inflationary multiverse truly accounts for fine-tuning, then (unless there are a rather large number of scalar fields) the inflationary epoch must go back much earlier than the current observational lower bound. This expectation contrasts sharply with the Hartle–Hawking wavefunction proposal [Hartle:1983ai]. Although the proposal appears to give reasonable predictions for the high-energy modes of the quantum fluctuations [Mukhanov:1981xt, Starobinsky:1982ee, Guth:1982ec, Hawking:1982cz, Bardeen:1983qw], it favors an inflationary period with far fewer e-folds than even observational bounds permit [Maldacena:2024uhs]. In fact, if we do not condition on inflation happening at all, the Hartle-Hawking proposal predicts that an even more dominant contribution to the saddle should be a big empty de Sitter universe with the present-day value of $\Lambda$, dramatically contradicting observation.
These issues could potentially be remedied by considering another variant of the no-boundary proposal, for example the Vilenkin tunnelling proposal [^19] This model was originally proposed in [vilenkin1982creation], for more recent discussion see Vilenkin and Yamada [Vilenkin:2018dch, Vilenkin:2018oja] and citations therein. in which the universe tunnels from zero size ($a=0$) along a Lorentzian contour. Such models typically make the opposite prediction from the Hartle-Hawking model, that the universe should start out as small as possible, and thus at a high value of the inflationary potential. It was argued in [Feldbrugge:2017kzv] that this model gives an unnormalizable (inverse Gaussian) prediction for the quantum fluctuations, but Vilenkin and Yamada claim that this issue can be resolved by adding an appropriate boundary term to the action; see also [DiTucci:2019dji] for a similar idea.
Thus, if the initial condition was not Hartle-Hawking, it is more plausible that the universe started at a point in the landscape from which a large amount of inflation was possible. It is also possible for inflation never to end (see e.g. [Guth:2007ng] for a review of this “eternal inflation”), but as such scenarios typically involve quantum fluctuations, the above conclusions will not necessarily be applicable unless the FEB can be generalized to quantum situations. This will be the subject of the next section.
@@ 5 Quantum effects
We now comment on a possible quantum generalization. As stated above Eq. (8), we assumed the NEC in this derivation. However, the NEC can be violated by quantum effects [Epstein:1965zza] such as Hawking radiation. We need to consider a quantum generalization of the NEC, which will involve accounting for entropy variation. We will keep Newton’s constant, $G_{\textrm{N}}$, explicit in this section because its renormalization will play an important role.
@@
associated with a given $\sigma_{\lambda}(y_{\perp})$. $S_{\text{gen}}$ generalizes the notion of area when quantum effects are important. Rather than studying how the entropy and area evolve as a function of $\lambda$, we treat the area and entropy as functionals of $\sigma_{\lambda}(y_{\perp})$ and consider functional derivatives localized in a small area $\mathcal{A}$ around $y_{1,\perp}$; see Fig. 3. Using the same notation as Ref. [Bousso:2015mna], we use the shorthand
@@
$\frac{d\theta}{d\lambda}\leq\frac{-1}{D-2}\theta^{2}\ ,$ (55)
@@
motivated by the idea that quantum focusing inequality should have a definite-sign quadratic term just as the classical Raychaudhuri Eq. does. Then we consider the contribution of a coherently varying field along the light ray. We refer to Eq. (56) as the SQFC, because the original quantum focusing conjecture (QFC) only imposed that the left-hand side was negative [Bousso:2015mna].
@@
In the original QFC, the entropy suffers from UV divergences, but these can be absorbed into the renormalization of Newton’s constant, so $\Theta$ is well-defined [Susskind:1994sm, Bousso:2015mna]. More concretely, the leading order area-law renormalization leads to the replacements
@@
for some constant $b$, where $G_{\textrm{N}}$ is Newton’s constant. The shift in Eq. (58) leaves $S_{\textrm{gen}}$ (52) invariant. In contrast, a shift in Newton’s constant due to renormalization changes the relative sizes of the two terms in Eq. (57) due to the explicit factor of $G_{\textrm{N}}$. The meaning of the bound thus requires picking some choice of $G_{\textrm{N}}$, which is a nontrivial decision if $G_{\textrm{N}}$ is subject to RG flow. As a conservative choice, we propose that whenever the SQFC is applied to an EFT calculation valid between some minimum energy scale $E_{\mathrm{min}}$ and maximum energy scale $E_{\mathrm{max}}$, that (i) an effective value of $G_{N}(E)$ should be defined at each energy $E_{\text{min}}<E<E_{\text{max}}$ using the area-law formula, and that (ii) the SQFC holds for whichever of these values of $G_{N}(E)$ is the smallest (i.e. the least stringent choice in (57)).
@@
and an “area law” scheme for $G_{\text{N}}$ will set $c_{1}=0$, so that the area-law part of the generalized entropy is entirely contained in the $A/4G_{\text{N}}$ term. A more complicated story occurs if the matter sector flows from a UV CFT to an IR CFT. In this case, Eq. (59) only holds asymptotically in either the $r\to 0$ limit (the UV CFT) or the $r\to\infty$ limit (the IR CFT). It has been proven that $S^{\prime\prime}(r)<0$ [Casini:2012ei], so it is not possible to have $c_{1}=0$ in both the UV and the IR. In fact, $G_{\text{UV}}<G_{\text{IR}}$ (i.e. in this area-law scheme gravity is anti-screening). In the case of the future lightcone of a point $p$ in the 2+1 Minkowski vacuum, we have checked that the SQFC holds when defined in terms of $G_{\text{UV}}$, by setting $c_{1}=0$ in the UV. (This implies that $S^{\prime}(r)\leq 0$ for all $r>0$, and thus ensures that $\theta$ and $S^{\prime}$ have the opposite signs.)
@@
In higher dimensions, it might be necessary to also think carefully about subleading divergences, beyond the area law, which are associated with higher curvature terms in the effective gravitational action (cf. [Bousso:2015mna] and citations therein).
@@
The strengthened QFC implies some desirable properties about spacetime, such as quantum generalizations of the Penrose singularity theorem [Penrose:1964wq, Hawking:1965mf, Hawking:1973uf]. It was previously argued by one of us that the Generalized Second Law (GSL), which is implied by the QFC, can be used to prove the incompleteness of null geodesics from the existence of a “quantum trapped surface” $\cal T$ with $\Theta\leq 0$ everywhere [Wall:2010jtc]. [^20] Together with the auxiliary assumptions of global hyperbolicity and spatial noncompactness, which are also used in the classical Penrose theorem. The SQFC would allow one to strengthen this result by deriving that the termination should occur before an affine time $\Delta\lambda\leq-(D-2)/\Theta[\mathcal{T}]$, analogously to the classical result.
@@
We further note that the SQFC is similar to a bound proposed by Ben-Dayan [Ben-Dayan:2023inz]. Although the details of Ben-Dayan’s inequality vary from our own, it is similarly obtained by adding to the QFC a term which vanishes when $\Theta=0$. All such bounds will imply the restricted QNEC, that $d\Theta/d\lambda\leq 0$ whenever $\Theta=0$, which has been proven in the context of the brane-worlds scenario [Shahbazi-Moghaddam:2022hbw].
@@
We now turn back to bounding field excursions. The core intuition behind the original FEB’s derivation from the NEC is that, if the other stress–energy contributions violated the NEC, one could simply eliminate the field excursion and reveal a NEC violation. We adopt a similar line of reasoning here. We assume that the expectation values of the scalar field and the area operator obey Einstein’s equations. To derive a quantum FEB (QFEB), we consider an original state where the scalar expectation value $\langle\phi\rangle$ changes along the null ray. We then act on this state with an operator that removes the coherent variation of $\phi$ within a local region and apply Eq. (56) to this state. Thus, in the new state $\langle\phi\rangle=\textrm{constant}$. (As the local region can be small, the required shift in $\Delta\langle\phi\rangle$ can be made arbitrarily small, and thus should not significantly affect the behavior of the matter quantum state through direct couplings.) In order for Eq. (56) to hold for the new state, the following bound must hold on the original state with coherent field variation: